A certain weak base has a Kb of 7.80  10-7. What concentration of this base will produce a pH of 10.16? ______M
Answer #1
Concepts and reason

The ratio of concentration of product to the concentration of reactant is represented as equilibrium constant. The equilibrium constant is expressed as.

Base dissociation constant represents the strength of a base in a solution. The strength of the base is represented by the equilibrium constant expression.

Fundamentals

The expression forcan be written and calculated as follows:

Example:

A+H,0 = AH*+OH
K, - [AH ][oH]
[A]

pH+pOH=14
pH=-log[H]
pOH =-log(OH)

Then

H* )=concentration of H*
[OH =concentration of OH®

ICE-table is the method of expressing the concentration of species which under go equilibrium.

I=Initial concentration
C=Change in concentration
E=Equilibrium concentration

B + H20
HB +OH
K. = [HB*](OH)
[B]

pH = 10.16
pOH = 3.84 (14-10.16)
[OH-] = 1038
[OH-] = 1.445*10*

ICE table can be written as

B+ H,0 = [BH] + [ou]
I(M) X
0 0
C(M) -1.445*10* +1.445*10* +1.445*10*
E(M) x-1.445*10* +1.445*10* +1.445*10*

Since equilibrium constant was given K= 7.80x107

[BH][on]
[B]
7.80x107- (1.445x104][1.445x1047
X - 1.445x104

Then on rearranging the above equation

[1.445*10*1.445x10^]
X-1.445x10=
7.80x107
x-1.445*10* = 0.02677
x = 0.02677 +1.445x104
[B] =0.02691M

Ans:

The concentration of base is given below:

[B]=0.02691M